Problem 2 (Ignore taxes for this problem)

Bullock Prosthetics is planning to buy 3-D printing machinery costing $380,000. This machinery’s expected useful life is 5 years. They require a minimum rate of return of 8%, and have calculated the following data pertaining to the purchase and operation of this machinery:

Year

Estimated Annual

Cash Inflows

Estimated Annual

Cash Outflows

Depreciation

1

$ 90,000

$15,000

$60,000

2

$150,000

$45,000

$60,000

3

$230,000

$95,000

$60,000

4

$270,000

$110,000

$60,000

5

$300,000

$125,000

$60,000

  1. Determine Terra's payback period, accounting rate of return, and NPV for this investment?
Answer #1

Working Note:-

Year Estimated Annual Estimated Annual Depreciation (c ) Net Income (a-b-c) Cummulative Cash Inflow
Cash Inflows (a) Cash Outflows (b)
1 $90,000 $15,000 $60,000 $15,000 $90,000
2 $150,000 $45,000 $60,000 $45,000 $240,000
3 $230,000 $95,000 $60,000 $75,000 $470,000
4 $270,000 $110,000 $60,000 $100,000 $740,000
5 $300,000 $125,000 $60,000 $115,000 $1,040,000
Total $1,040,000 $390,000 $300,000 $350,000
ARR= Average Income / Average Investment
=$70000/$190000= 46.67%
=Average Income= 350000/5= $70000
Average Investment =(Investment- Residual Value)/2
(380000-80000)/2=150000
Payback period = 2+ (380000-240000)/230000
=2.61 Year
Computation of NPV
Year Estimated Annual Estimated Annual PVAF @8% NPV
Cash Inflows (a) Cash Outflows (b) ( C) (a-b)Xc
Initial Investment $0.00 $380,000.00 1.0000 -$380,000.00
1 $90,000.00 $15,000.00 0.9259 $69,444.44
2 $150,000.00 $45,000.00 0.8573 $90,020.58
3 $230,000.00 $95,000.00 0.7938 $107,167.35
4 $270,000.00 $110,000.00 0.7350 $117,604.78
5 $300,000.00 $125,000.00 0.6806 $119,102.06
5 $80,000.00 0.6806 $54,446.66
Net Presetn Value $177,785.86